OpenStudy (anonymous):

Mary delivers 32 newspapers to her customers in 25 minutes. These deliveries represent 64% of her customers. How many customers does Mary have? A. 50 B. 57 C. 16 D. 20

OpenStudy (mathmale):

In this case, the "25 minutes" is irrelevant. 32 deliveries constitute 64% of Mary's customers. Let x=the total number of customers Mary serves. Can you write an equation in x, using the above info? Can you then solve it for x?

OpenStudy (anonymous):

32 is 64% of what number translate directly in to math 32 means 32 is means equals 64% means .64 of means times what number is what you are looking for, lets call it ? \[32=.64\times ?\]

OpenStudy (anonymous):

so what is the exact ans

OpenStudy (anonymous):

32/.64 = 50

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you

OpenStudy (mathmale):

7259: Both Satellite and I have tried to provide info to help you find your own answer. You may have to do this on a test or quiz, where you can't ask anyone for "the exact right answer." Would you please not ask for "answer" in the future?

OpenStudy (mathmale):

Instead, please ask for guidance towards finding the correct answer yourself.

OpenStudy (the_fizicx99):

32 / 0.64 = 50 So 50 customers, 50 - 32 = 18 100 - 64 = 36 18 = 0.36x /0.36 /0.36 So 50 is right

OpenStudy (the_fizicx99):

The other 36% is 18 btw >.<